OpenStudy (anonymous):

Calculate sin(arctan2-arcsin(1/sqrt(10))), (No calculator).

OpenStudy (anonymous):

I'm initially thinking of using sin(u-v)=sinucosv-sinvcosu, which leads me to: sin(arctan2)*cos(arcsin(1/sqrt10))-cos(arctan2)/sqrt10 Quite stuck here, is there any other path that is easier, is there something someone sees that I'm missing? :)

OpenStudy (agent0smith):

Draw some triangles 

OpenStudy (anonymous):

I don't follow. :(

OpenStudy (agent0smith):

let \[\Large \theta = \arcsin \frac{ 1 }{ \sqrt{10} }\] then make the triangle above

OpenStudy (anonymous):

I follow for that; I don't see how it applies to calculating our main thing though. :/

OpenStudy (anonymous):

\[\sin(\arctan(2)-\arcsin(\frac{1}{\sqrt{10}}))\]

OpenStudy (agent0smith):

It simplifies the arcsin, you can now find cosine. (use the same triangle) cos(arcsin(1/sqrt10)) can be simplified.

OpenStudy (anonymous):

How?

OpenStudy (agent0smith):

I think you were on the right track with this: sin(arctan2)*cos(arcsin(1/sqrt10))-cos(arctan2)/sqrt10

OpenStudy (agent0smith):



OpenStudy (agent0smith):

cos(arcsin(1/sqrt10)) remember that... \[\Large \theta = \arcsin \frac{ 1 }{ \sqrt{10} }\] so you want cos theta.

OpenStudy (agent0smith):

\[\Large \cos(\arcsin(1/\sqrt10)) = \cos \theta\] since theta is = arcsin(1/sqrt10)

OpenStudy (agent0smith):

You can do the same with arctan2. Make a triangle, find the other sides... let arctan2=phi, then you need sin phi, since you have a sin(arctan2)... sin(arctan2)*cos(arcsin(1/sqrt10))-cos(arctan2)/sqrt10

OpenStudy (agent0smith):

Oh and you need cos phi, since there's a cos(arctan2)

OpenStudy (anonymous):

I see! So you're going back to the triangle after you've switched them in?

OpenStudy (anonymous):

I think I understand what you mean :)

OpenStudy (agent0smith):

Yes, that's really the only way to do these... make triangles to get rid of inverses... like sin(arccosx) can be simplified because arccosx = some angle. So it's sin(some angle). The triangles just help you find all the other sides you need for tan, sin, cos of that angle. Or secant, cosecant, cotangent too.

OpenStudy (anonymous):

We never utilize secants etc in Swedish math I think. Really like your way of solving this! Man I love these calculations, so simple yet so deep! :D

OpenStudy (anonymous):

(In our undergrad courses that is)

OpenStudy (agent0smith):

haha yes these are fun. But they are pretty confusing at first.. and even sometimes later, i'll think "wait, what the hell was i doing again...?"

OpenStudy (anonymous):

Indeed! Haha

OpenStudy (agent0smith):

But if you simplify them bit by bit they're easier, and make sure to name your angles... eg let phi =arctan2 and so on. That way you aren't remember what's what in your head.

OpenStudy (anonymous):

Did you solve for the answer? It's beautiful as well! :D

OpenStudy (agent0smith):

looks like sin(arctan2-arcsin(1/sqrt(10))) = (sqrt2)/2 which means... arctan2-arcsin(1/sqrt(10)) = 45 degrees.

OpenStudy (anonymous):

;)